# Interpreting the Spectro-Temporal Properties of the Black Hole Candidate Swift J151857.0-572147 during its First Outburst in 2024

### v1

@@ 3.1.1 Outburst Profile, and Hardness Ratio
The BHC Swift J151857.0-572147 was not observed by the MAXI/GSC instrument. The source is located at $\sim 0.2^{\circ}$ from the source Cir X-1. Although the facility could identify the brightening of the source, the two sources could not be resolved seperately. In Figure 1, we show the location of the two sources in the upper panel. It can be noticed that the two sources are located very close to each other. The lower panel of the figure shows increased activity due to the outburst of Swift J151857.0-572147.
@@
In Figure 2, we show the variation of the count rates for around 15 days. The count rates are extracted using LE, ME, and HE light curves of HXMT in the $2-10$, $10-35$, and $27-250$ keV energy bands. In panel (a), we show the variation of those source and background count rates for the three bands (in respective colors). Red is for LE, while green and blue colors are used to represent ME and HE bands. The filled circle (of each color) lines represent the source counts, whereas the triangle-shaped lines represent the background count rates. As can be noticed, the HE background count rate was quite high and was almost comparable to the source count rate. The other two bands showed a significant difference in count levels between source and background. In Table 2, we list the start, end, and average MJDs of all our analyzed exposures. We also list the source and background count rates for LE, ME, and HE in Table 2. In panel (b) of Figure 2, we show the count rates in $2-4$, $4-10$ keV energy bands, which are extracted using LE light curves. In panel (c), the hardness ratio (HR) is plotted using the ratio of the LE count rates of $4-10$ to $2-4$ keV.
From the light curves, we see that the source is quite bright during the outburst. The LE count rate shows a smoother variation than the ME and HE bands, which can also be noticed in the (b) panel. From the variation, HR gives a rough idea that the source had already moved past its hard state as Insight-HXMT started monitoring the source. As time progressed, spectral nature progressed from intermediate to softer states. However, we need timing and spectral analysis results to designate this firmly. We discuss them in the next two subsections.
@@ 3.1.2 Low-Frequency Quasi Periodic Oscillations
Figure 3: Model fitted power density spectrum (PDS) continuum in $0.2-50~{}Hz$ for (a) LE, (b) ME, and (c) HE. The best fit is achieved using a combination of a set of models: power-law, linear, constant, and Lorentzian.
@@
We created the PDS to analyze QPOs using the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE), as stated in §2. In Figure 3, we show the best model-fitted PDS continuum for the three bands (a) LE, (b) ME, and (c) HE for the observation ID P0614374001 (exposure ID: P061437400101-20240304-01-01). While, both the QPO and harmonic were present in the LE band, the harmonic nature was absent in ME and was not very prominent in HE. The QPO and harmonic have a 1:2 ratio in frequency with the $\nu_{harmonic}\sim 6.43\pm 0.04$ Hz. The harmonic in this exposure has an FWHM of $0.39\pm 0.13$ and normalization of $0.57\pm 0.14$. We discover that each of the three energy bands’ light curves has a fundamental QPO nature. We first checked all the exposures for the observation ID P0614374001. From our fitting, we first extracted the basic QPO information, which are QPO frequency ($\nu_{qpo})$, full-width at half-maximum (FWHM or LW), and QPO normalization (LN).
@@
We found that QPO was present in most of the exposures of this observation ID. It was present for all exposures in the ME band and was absent in the last LE band and second and fifth HE band. Also, the QPO frequency evolved within the short period of the duration of this observation ID. Thus, we checked for QPOs for all the exposures. At the start of our analysis period, fundamental QPO was present in almost most of the exposures. The $\nu_{qpo}$ was $\sim$ 3.2 for all three bands on MJD 60373.9, and it increased as the outburst progressed. Then after some days, it decreased, and then again showed an increasing trend. Then, it again decreased and increased and decreased and continued this way. The highest frequency in the LE band was 8.1 Hz on MJD 60376.9, on which both the light curves in the ME and HE bands were not created by the hpipeline command. The highest frequency in the ME band was 8.97 Hz on MJD 60377.9, on which the LE and ME light curves were not produced. In the HE band, $\nu_{qpo}$ was the highest on MJD 60379.3 with a value of 6.82 Hz. We show the variations of the QPO frequency during our full analysis period in Figure 4(a-c) for (a) LE, (b) ME, and (c) HE. In Table 3, we listed the values of $\nu_{qpo}$ in columns 2, 3, & 4 for LE, ME, and HE. Although for the exposure P061437400103-20240305-02-01, there was a presence of a harmonic nature in the HE band, we did not fit it as the noise was high and the harmonic was like a broad Lorentzian feature. We did not find harmonic for any other exposures of any other observation ID in any of the three bands.
@@
We were able to extract certain information about QPOs, such as full-width at half maximum (FWHM) and Normalization (LN) by the use of PDS fitting. For the exposures, we additionally retrieved the source and background count rates. Using the formula from Bu et al. (2015), we estimated the fractional RMS as $RMS=\sqrt{\frac{P}{S+B}}\frac{S+B}{S}$, which denotes the fractional variability in the PDS. Here, $S$, and $B$ represent the count rates of the source and the background, respectively. $P$ is the Leahy normalized power. We also estimated the $Q$-factor ($\nu/\delta\nu$), which measures the sharpness of the QPO. Table 3 lists these values for LE, ME, and HE in columns 5–7 ($Q$-value) and 8–10 (RMS), respectively. This is shown in Figure 5. The variations of the $Q$ factor were consistent in the three different bands. To check if there is any correlation between the QPO RMS and QPO frequency, we plotted those two properties against each other in Figure 6. We have not found any correlation between them for this source in all three energy bands.
As explained before, we also checked the energy dependence of QPOs using the HE light curve in 7 different energy bands. These energy ranges were chosen to maintain similarity with Ma et al. (2023). The PDS continuums for the observation ID P0614374001 (exposure ID: P061437400101-20240304-01-01) are given in Figure 7(a-g) for respective energy bands. For this exposure, we find that the fundamental QPO was prominently present at 3.274 Hz in the $27-35$ keV energy band, while it is also present in the $35-48$ keV with a little change of frequency of 3.222 Hz. However, the nature of QPO was not as strong as in the $27-35$ keV. Above 48 keV, we did not find any nature of fundamental QPO. We notice a sharp fall of QPO strength above 48 keV. Chatterjee et al. (2021) studied QPO energy dependence for the BHC GRS 1716-249 using AstroSat data. Although the fundamental QPO nature got weaker in high energies in that report, it did not show this type of sharp fall of QPO nature after some energy band. A possible weak harmonic nature was noticed in the $35-48$ keV band, which was not present in the $27-35$ keV. However, it looks very weak and we did not model it. Harmonic nature was also not observed above 48 keV. We checked this for all the 31 exposures for which fundamental LFQPO was present in the HE light curve. We find that for all of these exposures, QPO was absent above 48 keV. For some exposures, we found that LFQPO was absent in the 35-48 keV band, although it was present in the 27-35 keV band.
@@
Using the formulas, as mentioned above, we estimated $Q$-values and RMS (%) for all these exposures in both these bands. In Figure 8, we show the variation of QPO frequency, $Q$-value, and RMS (%) for all these exposures with time. We notice that the $\nu_{qpo}$ varied in a very narrow range between these two energy bands, which is within the error range. The $Q$-value shows a random variation for both the bands, where it was sometimes higher for 27-35 keV bands and sometimes for 35-48 keV. The overall variation of RMS (%) was higher in case of 27-35 keV band, compared to the higher band. The values of the variation of QPO properties in case of energy dependence is given in the Table 4.
@@
We also show the variation of QPO RMS with energy in Figure 9. We noticed that the QPO RMS was lowest in the LE band. It was the highest in the ME band. Then it started to decrease. Above 48 keV, we did not find the presence of any QPO.
@@
Apart from this, we also searched for QPOs in the $48-250$ keV energy band light curve. Since above $48$ keV, no QPO was found, we wanted to check if the energy range was higher, it could show QPO nature or not. In all of the exposures, except one, we did not find the presence of QPO in this energy band. This is for the observation ID. P0614374001 (exposure ID. P061437400103-20240305-02-01). We found that there was the presence of a fundamental QPO in this exposure at $3.06\pm 0.05$ Hz. This is shown in the Figure 10.
@@ 3.1.3 High Frequency QPOs
Apart from looking for low-frequency QPOs, we also searched for high-frequency QPOs (HFQPOs) in all the light curves for the three bands in all 62 exposures. In Figure 11(a-c), we show the PDS continuum for $0.001$ sec time-binned (Nyquist frequency = 500 Hz) curve for (a) LE, (b) ME, and (c) HE. However, we did not find any signature of HFQPOs in any of our light curves. The frequency in the PDS in LE, ME, and HE in Figure 11, are similar to those in Figure 3. Those are the LFQPOs present in those light curves during that exposure.
@@ 3.2 Spectral Properties
Studying the spectrum features sheds additional light on the nature of the outburst in addition to the temporal properties. We examined the source using the Insight-HXMT data that was accessible, for 14 exposures in total. The exposure IDs in Table 2’s first column have a ‘*’ symbol next to them. We perform a thorough spectral study using HXMT data on this source for every consecutive days for the available data. Our spectrum investigation was initiated with MJD 60373.9. For spectral fitting, we have simultaneously analyzed LE+ME+HE in the $2-100$ keV energy band (LE in 2–10, ME in 10–35, and HE in 27–100 keV) for all of the chosen exposure IDs.
Model Fitted Insight-HXMT Spectra
Refer to captionRefer to captionRefer to caption
@@
Figure 12: Best model fitted unfolded spectra for observation ID. P0614374001 (Exposure: P061437400101-20240304-01-01) using (a) Model-1, (b) Model-2, and (c) Model-3.
@@
First, we tried to model the spectrum with simple additive models diskbb and power-law. We also used the multiplicative tbabs (with wilm abundance, Wilms et al. (2000)) model to account for the interstellar absorption. The model fitted unfolded spectrum is given in the Appendix section in Figure A.15. Although, the $\chi^{2}/DOF$ value was acceptable, we noticed that the spectrum changes its slope above $\sim 20$ keV. Thus, we replaced the power-law with the broken power-law model, which accounts for the change of slope after certain energy, called break energy ($E_{b}$). In XSPEC, our model was expressed as constant*tbabs(diskbb + broken powerlaw). The three distinct instruments (LE, ME, and HE) are normalized using the constant. This model could fit the spectra for an acceptable $\chi^{2}/DOF$ value ($\sim 1$). These values can be found in column 11 of Table 5. Although this model fit was acceptable, there was a reflection nature in the spectrum. To account for that, we replaced the broken power-law model with the reflection model in neutral medium pexrav. With this model, using the model combination as constant*tbabs(diskbb + pexrav), we also achieved the best fit. The $\chi^{2}/DOF$ values can be found in column 11 of Table 6. We also checked the reflection component by using the reflection model pexriv which takes ionization into account. Thus constant*tbabs(diskbb + pexriv) is our Model-3 combination in XSPEC. We also achieved best fit using this combination. Here, we like to point out our approach using the Model 3. Except for 2 parameters, all the parameters of this model are the same as the pexrav model. While fitting with this model, we fixed the cut-off energy ($E_{cut}$) of this model to the pexrav model. Also, we found while fitting that the disk temperature parameter (in units of Kelvin) was always taking its highest value, which is $10^{6}$ Kelvin. Thus, for all the spectral fitting using this model, we freeze the value of this parameter to this highest value. The extra parameter that this model has over pexrav is the disk ionization parameter ($\xi$), which is given as $\xi=4\pi F_{ion}/n$, where $n$ is the density of the reflector (Done et al. 1992) and $F_{ion}$ is the irradiating flux in the 5 eV to 20 keV energy band. The best fitted parameters are given in Table 7 along with the fitting statistics. We find that the parameter variation of the pexriv model is similar to the pexrav model. In Figure 12(a-c), we show the model fitted best unfolded spectra using (a) Model-1, (b) Model-2, and (c) Model-3.
@@
Figure 13: Variation of spectral model fitted properties. (a), (b), and (c) show the variation of hydrogen column density (in 1022 cm-2 unit), inner-disk temperature ($T_{in}$ in keV), and photon index ($\Gamma$) for all the three models with time. The red, blue, and green colors represent the properties for the Model-1, Model-2, and Model-3, respectively. In the (c) panel, we show the $\Gamma 1$ and $\Gamma 2$ of the Model-1 using red color filled and hollow squares, respectively.
In Figure 13, we show the variations of some of the spectrally analyzed properties from both models. In panel (a), we show the variations of the hydrogen column density ($N_{H}$) for both the model fittings (red filled-square for Model-1 and blue filled-square for Model-2). We notice that they show close variations within the error range throughout. While for Model-1, $N_{H}$ varied between $(4.3-6.5)\times 10^{22}$ cm-2, it varied in the range of $(5.1-6.3)\times 10^{22}$ cm-2 for Model-2 and in the range of $(5-6.9)\times 10^{22}$ for Model-3. In panel (b), we show the variations in the inner-disk temperature ($T_{in}$ in keV) for both models. $T_{in}$ shows variation in the range of $0.95-1.6$ keV for Model-1, $1.3-2.1$ keV for Model-2, and $0.9-2.1$ keV for Model-3. Since, the last exposure could not be fitted using Model-2, we think we have got the narrower variation using Model-2. In panel (c), we show the variations of photon indices. The red filled-square and hollow square represent $\Gamma 1$ and $\Gamma 2$ for the broken power-law model, where the blue and filled squares represent the $\Gamma$ of the pexrav and the pexriv models. The energy break ($E_{b}$) of the broken power-law model was $19.2\pm 0.3$ at the starting day and it decreased to $8.54\pm 0.3$ at the end of the analysis period. All the spectral parameters are given in Table 5, including fitting constants, needed to normalize simultaneous data in three bands. Similarly, all the spectral parameters are listed in Table 6 and 7, for the analysis using Model-2 and Model-3. For analysis with Model-2 and Model-3, we fixed the abundances to solar abundance and also varied the value of the inclination to a narrow range around $30^{\circ}$ as reported by Mondal et al. (2024). However, on the last date of our analysis period, we found that although the fitting statistics were acceptable, the parameter space was unphysical. We tried to adjust the parameter space. However, the best fit was not achieved then. Thus we have left it blank in Table 6.

### v2

@@ 3.1.1 Outburst Profile, and Hardness Ratio
Although the BHC Swift J151857.0-572147 was observed by the MAXI/GSC instrument, it was not recognized as a new source by them. The source is located at $\sim 0.2^{\circ}$ from the source Cir X-1. Although the facility could identify the brightening of the source, the two sources could not be resolved seperately. In Fig. 1, we show the location of the two sources in the upper panel. It can be noticed that the two sources are located very close to each other. The lower panel of the figure shows increased activity due to the outburst of Swift J151857.0-572147. According to Peng et al. (2024), Cir X-1 was active in the soft state during this time. Thus, the contribution from Cir X-1 could have contribution to the low energy part (i.e., LE) of the light curve.
@@
In Fig. 2, we show the variation of the count rates for around 15 days. The count rates are extracted using LE, ME, and HE light curves of HXMT in the $2-10$, $10-35$, and $27-250$ keV energy bands. In panel (a), we show the variation of those source and background count rates for the three bands (in respective colors). Red is for LE, while green and blue colors are used to represent ME and HE bands. The filled circle (of each color) lines represent the source counts, whereas the triangle-shaped lines represent the background count rates. As can be noticed, the HE background count rate was quite high and was almost comparable to the source count rate. This could be due to the combination of two reasons e.g., the large effective area of HE and the close proximity of the source to Cir X-1. However, as mentioned earlier, the contribution from Cir X-1 could be in low energy as it was in the soft state. Thus, it is hard to comprehend the actual reason behind it. The other two bands showed a significant difference in count levels between source and background. In Table 2, we list the start, end, and average MJDs of all our analyzed exposures. We also list the source and background count rates for LE, ME, and HE in Table 2. In panel (b) of Fig. 2, we show the count rates in $2-4$, $4-10$ keV energy bands, which are extracted using LE light curves. In panel (c), the HR is plotted using the ratio of the LE count rates of $4-10$ to $2-4$ keV.
From the light curves, we see that the source has high count rates in all three bands. For Insight-HXMT, a count rate of 800, 500, and 800 counts/sec from LE, ME, and HE corresponds to the flux equivalent to 1Crab. These flux values give the idea about the brightness of the source. From the variation, HR gives a rough idea that the source had already moved past its hard state as Insight-HXMT started monitoring the source. As time progressed, spectral nature progressed from intermediate to softer states. However, we need timing and spectral analysis results to designate this firmly. We discuss them in the next two subsections.
@@ 3.1.2 Low-Frequency Quasi Periodic Oscillations
Figure 3: Model fitted power density spectrum (PDS) continuum in $0.01-50~{}Hz$ for the LE band for fitting in (a) powspec and (b) xspec. The best fit is achieved using a combination of a set of models: power-law, constant, and Lorentzian.
@@
We have created the PDS to analyze QPOs using the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE), as stated in §2. In Fig. 3, we show the best model-fitted PDS continuum for the LE band, for both fittings in (a) powspec, and (b) xspec for the observation ID P0614374001 (exposure ID: P061437400101-20240304-01-01). First, we used the combination of power-law, constant, and Lorentzian models in powspec. After the best model fitting, the same is followed in xspec. From the xspec fitting panel, one can also observe the goodness of fitting, as well as the contributions from various fearues of the PDS, including the fundamental QPO and harmonic. While, both the QPO and harmonic were present in the LE band, the harmonic nature was absent in ME and was not very prominent in HE, as observed from the fittings in all the bands. The QPO and harmonic have a 1:2 ratio in frequency with the $\nu_{qpo}\sim 3.19\pm 0.02$ and $\nu_{harmonic}\sim 6.43\pm 0.04$ Hz, respectively. The fundamental and harmonic QPO in this exposure has an FWHM of $0.35\pm 0.05$, $0.39\pm 0.13$ and normalization of $1.84\pm 0.23$, $0.57\pm 0.14$, respectively. We noticed the presence of fundamental QPOs in the PDS of each of the three energy bands. We first checked all the exposures for the observation ID P0614374001. From our fitting, we first extracted the basic QPO information, which is QPO frequency ($\nu_{qpo})$, full-width at half-maximum (FWHM or LW), and QPO normalization (LN).
@@
We found that QPO was present in most of the exposures of this observation ID. It was present for all exposures in the ME band and was absent in the last LE band and second and fifth HE band. Also, the QPO frequency evolved within a short period of $\sim 1.5$ days of the duration of this observation ID P0614374001. Thus, we checked for QPOs for all the exposures. At the start of our analysis period, fundamental QPO was present in almost all of the exposures. The $\nu_{qpo}$ was $\sim$ 3.2 for all three bands on MJD 60373.9, and it increased as the outburst progressed. Then after some days, it decreased, and then again showed an increasing trend. Then, it again decreased and increased and decreased and continued this way. The highest frequency in the LE band was 8.1 Hz on MJD 60376.9, on which both the light curves in the ME and HE bands were not created by the hpipeline command. The highest frequency in the ME band was 8.97 Hz on MJD 60377.9, on which the LE and ME light curves were not produced. In the HE band, $\nu_{qpo}$ was the highest on MJD 60379.3 with a value of 6.82 Hz. We show the variations of the QPO frequency during our full analysis period in Fig. 4(a-c) for (a) LE, (b) ME, and (c) HE. In Table 3, we listed the values of $\nu_{qpo}$ in columns 2, 3, & 4 for LE, ME, and HE. Although for the exposure P061437400103-20240305-02-01, there was a presence of a harmonic nature in the HE band, we did not fit it as the noise was high and the harmonic was like a broad Lorentzian feature. We did not find harmonic for any other exposures of any other observation ID in any of the three bands.
@@
We were able to extract certain information about QPOs, such as full-width at half maximum (FWHM) and Normalization (LN) by the use of PDS fitting. For the exposures, we additionally retrieved the source and background count rates. Using the formula from Bu et al. (2015), we estimated the fractional RMS as $RMS=\sqrt{\frac{P}{S+B}}\frac{S+B}{S}$, which denotes the fractional variability in the PDS. Here, $S$, and $B$ represent the count rates of the source and the background, respectively. $P$ is the Leahy normalized power. We also estimated the $Q$-factor ($\nu/\delta\nu$), which measures the sharpness of the QPO. Table 3 lists these values for LE, ME, and HE in columns 5–7 ($Q$-value) and 8–10 (RMS), respectively. This is shown in Fig. 5. The variations of the $Q$ factor were consistent in the three different bands. To check if there is any correlation between the QPO RMS and QPO frequency, we plotted those two properties against each other in Fig. 6. We have not found any correlation between them for this source in all three energy bands.
@@
As explained before, we also checked the energy dependence of QPOs using the HE light curve in 7 different energy bands. These energy ranges were chosen to maintain similarity with Ma et al. (2023). The PDS continuums for the observation ID P0614374001 (exposure ID: P061437400101-20240304-01-01) are given in Fig. 7(a-g) for respective energy bands. For this exposure, we find that the fundamental QPO was prominently present at 3.274 Hz in the $27-35$ keV energy band, while it is also present in the $35-48$ keV with a little change of frequency of 3.222 Hz. However, the nature of QPO was not as strong as in the $27-35$ keV. Above 48 keV, we did not find any nature of fundamental QPO. We notice a sharp fall of QPO strength above 48 keV. Chatterjee et al. (2021) studied QPO energy dependence for the BHC GRS 1716-249 using AstroSat data. Although the fundamental QPO nature got weaker in high energies in that report, it did not show this type of sharp fall of QPO nature after some energy band. A possible weak harmonic nature was noticed in the $35-48$ keV band, which was not present in the $27-35$ keV. However, it looks very weak and we did not model it. Harmonic nature was also not observed above 48 keV. We checked this for all the 31 exposures for which fundamental LFQPO was present in the HE light curve. We find that for all of these exposures, QPO was absent above 48 keV. For some exposures, we found that LFQPO was absent in the 35-48 keV band, although it was present in the 27-35 keV band.
Using the formulas, as mentioned above, we estimated $Q$-values and RMS (%) for all these exposures in both these bands. In Fig. 8, we show the variation of QPO frequency, $Q$-value, and RMS (%) for all these exposures with time. We notice that the $\nu_{qpo}$ varied in a very narrow range between these two energy bands, which is within the error range. The $Q$-value shows a random variation for both the bands, where it was sometimes higher for 27-35 keV bands and sometimes for 35-48 keV. The overall variation of RMS (%) was higher in case of 27-35 keV band, compared to the higher band. The values of the variation of QPO properties in case of energy dependence is given in the Table 4.
@@
We also show the variation of QPO RMS with energy in Fig. 9. We noticed that the QPO RMS was lowest in the LE band. It was the highest in the ME band. Then it started to decrease. Above 48 keV, we did not find the presence of any QPO.
@@
Apart from this, we also searched for QPOs in the $48-250$ keV energy band light curve. As we move to the higher energies, the detector’s effective area decreases. Thus, the number of photons also reduces. However, in high energies (e.g., $\sim 50$ keV), if the energy range is large, the PDS of the light curve may show the presence of QPOs. This is what we wanted to check. Since above $48$ keV, no QPO was found in the PDS, we wanted to make a further consistency check in a broader energy band, if it shows any QPOs or not. We could not find the presence of QPO in the higher energy bands in any of the exposures, except for the observation ID. P0614374001 with exposure ID. P061437400103-20240305-02-01. We found that there was the presence of a fundamental QPO in this exposure at $3.06\pm 0.05$ Hz. This is shown in Fig. 10. For all the detected QPOs, we have estimated the significance ($\sigma$) using the relations given in Sreehari et al. (2019). The values of the significance are given in each table where QPO information is listed.
@@ 3.1.3 High Frequency QPOs
Apart from looking for low-frequency QPOs, we also searched for high-frequency QPOs (HFQPOs) in all the light curves for all three bands in all 62 exposures. In Fig. 11(a-c), we show the PDS continuum for $0.001$ sec time-binned (Nyquist frequency = 500 Hz) curve for (a) LE, (b) ME, and (c) HE. However, we did not find any signature of HFQPOs in any of our light curves. The frequency in the PDS in LE, ME, and HE in Fig. 11, are similar to those in Fig. 3. Those are the LFQPOs present in those light curves during that exposure.
@@ 3.2 Spectral Properties
Studying the spectral features sheds additional light on the nature of the outburst in addition to the temporal properties. We examined the source using the Insight-HXMT data that was available for 14 exposures in total. The exposure IDs in Table 2’s first column have a ‘*’ symbol next to them. We perform a thorough spectral study using HXMT data on this source for every consecutive days for the available data. Our spectrum investigation was initiated with MJD 60373.9. For spectral fitting, we have simultaneously analyzed LE+ME+HE in the $2-100$ keV energy band (LE in 2–10, ME in 10–35, and HE in 27–100 keV) for all of the chosen exposure IDs.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 12: Best model fitted unfolded spectra for observation ID. P0614374001 (Exposure: P061437400101-20240304-01-01) using (a) Model-1, (b) Model-2, (c) Model-3, and (d) Model-4.
@@
First, we tried to model the spectrum with simple additive models diskbb and power-law. We also used the multiplicative tbabs (with wilm abundance, Wilms et al. 2000) model to account for the interstellar absorption. The model fitted unfolded spectrum is given in the Appendix section in Fig. 15. Although the $\chi^{2}/DOF$ value was acceptable, we noticed that the spectrum changes its slope above $\sim 20$ keV. Thus, we replaced the power-law with the broken power-law model, which accounts for the change of slope after certain energy, called break energy ($E_{b}$). We call this model as Model-1. The three distinct instruments (LE, ME, and HE) are normalized using the constant. Although this model fit was acceptable, there was a reflection nature in the spectrum. To account for that, we replaced the broken power-law model with the reflection model in neutral medium pexrav (Model-2). With this Model-2, we also achieved the best-fit statistics. Then we checked the reflection component by using the reflection model pexriv (Model-3) which takes ionization into account. Using the Model-3, as mentioned before, we achieved the best fit statistics. Here, we like to point out our approach using the Model-3. Except for 2 parameters, all the parameters of this model are the same as the pexrav model. While fitting with this model, we set the cut-off energy value to the one obtained from the fit with the pexrav model. Also, we found while fitting that the disk temperature parameter (in units of Kelvin) was always pegging to the highest value of $10^{6}$ Kelvin. Thus, for all the spectral fitting using this model, we freeze the value of this parameter to this highest value. The extra parameter that this model has over pexrav is the disk ionization parameter ($\xi$), which is given as $\xi=4\pi F_{ion}/n$, where $n$ is the density of the reflector (Done et al. 1992) and $F_{ion}$ is the irradiating flux in the 5 eV to 20 keV energy band. For the analysis with Model-2 and Model-3, we fixed the abundances to solar abundance and also varied the value of the inclination to a narrow range of around $30^{\circ}$ as reported by Mondal et al. (2024).
For all the fitting processes mentioned so far, the results are achieved by keeping the $N_{H}$ free, and it varies in a broad range. Therefore, we take the average of $N_{H}$ from all three models which is $\sim 5.6\times 10^{22}cm^{-2}$. Then we reanalyzed all the observations using the same models (Model-1 to 3) combinations by keeping $N_{H}$ fixed to the average value. The parameters of the fit do not change significantly by keeping $N_{H}$ fixed to the average value. We report both the results by keeping $N_{H}$ fixed and free in the next paragraph. All the best-fitted model parameters and statistics are given in Tables 5 to 10.
Since none of these models takes into account the relativistic effects, we performed the same spectral analysis using the relativistic reflection model, relxill to see the relativistic effects. We consider this as Model-4, as mentioned before. For this model fitting, we used the average $N_{H}$ value and froze it to $5.6\times 10^{22}~{}cm^{-2}$ for all observations. The relxill model has a total of 14 parameters, of which several can be fixed to reasonable values in order to avoid degeneracy in the best fit parameters (Peng et al. 2024; Mondal et al. 2024). From this model fitting, we can get some valuable information about the source like its spin and disk inclination. To achieve the best fitting with Model-4, we fixed some of the model parameters. We fixed $R_{in}=1$ and $R_{out}=1000$, $index1=index2=3$, $R_{br}=15$. For being a Galactic source, the redshift ($z$) was always kept to 0. The best-fitted parameters and statistics for this model fitting are listed in Table 11. In Fig. 12(a-d), we show the model-fitted unfolded spectra using (a) Model-1, (b) Model-2, (c) Model-3, and (d) Model-4.
Refer to captionRefer to caption
@@
Figure 13: Best-fitted spectral model parameters are shown with time MJD. The panels (a), (b), and (c) show the variation of hydrogen column density (in 1022 cm-2 unit), inner-disk temperature ($T_{in}$ in keV), and photon index ($\Gamma$) for all the three models. The red, blue, and green colors represent the parameters for the Model-1, Model-2, and Model-3, respectively. In the panel (c), we show the $\Gamma 1$ and $\Gamma 2$ of the Model-1 using red color filled and empty squares, respectively. The panels (d) and (e) represent the same parameters as in panels (b) and (c) but for the $N_{H}$ fixed to $5.6\times 10^{22}$ cm-2.
In Fig. 13, we show the variations of some of the spectral parameters for all four models. In the top figure, we show the variations of the parameters when the $N_{H}$ was free, whereas, in the bottom figure, we show the variations of the same parameters when $N_{H}$ was fixed to an average value. In panel (a), we show the variations of the $N_{H}$ for three different models (red filled-square for Model-1, blue filled-square for Model-2, and green filled-square for Model-3). We notice that they show consistent variations within the error range throughout. The $N_{H}$ for Model-1 varies between $(4.3-6.5)\times 10^{22}$ cm-2, $(5.1-6.3)\times 10^{22}$ cm-2 for Model-2 and $(5-6.9)\times 10^{22}$ for Model-3. In panel (d), we show the variations in the inner-disk temperature ($T_{in}$ in keV) for all four models for a fixed $N_{H}$. The $T_{in}$ shows variation in the range of $0.9-1.9$ keV for all four models. In panel (e), we show the variations of photon indices. The red filled-squares and empty squares represent $\Gamma 1$ and $\Gamma 2$ for the broken power-law model, where the blue, green, and black filled squares represent the $\Gamma$ of the pexrav, pexriv, and relxill models. We notice that the photon index was high if we take into the consideration of the presence of type-C QPOs. We also note that both the $T_{in}$ and $Norm_{diskbb}$ values did not show typical variations as was observed for other BHCs (Remillard & McClintock 2006). When an outburst progresses, the normalization generally decreases from a high value. In this case, we have not seen that trend, rather it varied randomly. In general, $T_{in}$ increases as an outburst progresses, reflecting the movement of the disk inwards. Here, for all the model combinations, we did not notice that increasing profile. The pexriv model fitted ionization parameter is low here in the range of $2.4\times 10^{-13}-3.8\times 10^{-8}$, the reason of which is not clear. Since the ionization parameter is low, the estimated irradiation flux ($F_{ion}$) is also very low, which can be estimated using the relation $\xi=4\pi F_{ion}/n$, where $n$ is the density of the reflector (Done et al. 1992) and $F_{ion}$ is the irradiating flux in the 5 eV to 20 keV energy band.
The relxill model fits provided a broad range for both the spin and inclination parameters of the source to be $0.5-0.81$ and $10-41$ (in degrees). This also could be due to the degeneracy in model parameters and significant changes in $Norm_{diskbb}$. Specifically on the last exposure Id, we found that the $Norm_{diskbb}$ increased to a very high value, compared to other Ids and disk temperature decreased to 0.9, whereas the photon index increased to 3.3. However, the ranges of the spin and inclination parameters cover the estimations in Mondal et al. (2024). Since the last exposure Id returned a very high photon index, we conclude the spectral state as SS, as was reported earlier by Mondal et al. (2024). The absence of QPO also supports the same spectral state. We found that the ionization parameter in relxill model fitting was high ($>3$) and the iron abundance ($AFe$) varied from $1.1-4.9$. The reflection fraction varied in a range of $0.12-2.29$. The $E_{cut}$ also showed random variations. We discuss the possible reasons in the next section.
3.3 Evolution of the Shock
In the TCAF scenario, the oscillation of the shock produces QPOs. According to this model, the QPO frequency can be written as (Molteni et al. 1996, Chakrabarti et al. 2005),
[…]
where the following are represented, respectively: $c$, $G$, $M_{BH}$, $X_{s}$, and $R$; these are the speed of light, the gravitational constant, the mass of the BH, the shock location, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$).
Figure 14: Evolution of the location of the shock, estimated from observed $\nu_{qpo}$. The panel (a) shows the variation of $\nu_{qpo}$ with time and (b) shows the variation of $X_{s}$ with time.
The QPO frequencies ($\nu_{qpo}$) are estimated from timing analysis as discussed earlier and the shock locations during the outburst are estimated using the above relation in Eq. 1. We found that at the beginning of the outburst, the shock was located at a distance of $\sim 100~{}r_{s}$ from the BH (see Fig. 14b). Later, as the $\nu_{qpo}$ increased, the shock moved inwards, decreasing the size of the corona. After a few days, the shock became stable. Table 3 provides the values for the shock locations (columns 11–13).